What mass of sulphuric acid contains 1.4 mol of Oxygen
1
Expert's answer
2020-01-06T07:46:09-0500

Task:

What mass of sulphuric acid contains 1.4 mol of Oxygen?


Solution:

1) H2SO4 --> 4O

n(H2SO4) = n(O)/4;

n(H2SO4) = 1.4 mol / 4 = 0.35 mol;


2) n = m / M;

m = n * M

M(H2SO4) = 2*Ar(H) + Ar(S) + 4*Ar(O) = 2*1 + 32 + 4*16 = 98 (g/mol).

m(H2SO4) = n(H2SO4) * M(H2SO4) = 0.35 mol * 98 g/mol = 34.3 g.

m(H2SO4) = 34.3 g.


Answer: 34.3 grams of sulphuric acid contains 1.4 mol of Oxygen.